# Direct imaging of hydrogen-driven dislocation and strain field evolution in a stainless steel grain

### v1

@@ Abstract
David Yang Mujan Seif Guanze He Kay Song Adrien Morez Benjamin de Jager Ross J. Harder Wonsuk Cha Edmund Tarleton Ian K. Robinson Felix Hofmann
@@ 1 Introduction
Hydrogen is an attractive energy carrier that can be produced through water electrolysis using renewable resources [1]. This green hydrogen can be used to decarbonize industries that rely heavily on fossil fuels such as aviation [2], heavy transportation [3], industrial production [4], and power generation [1]. As countries strive to reach their Net Zero 2050 sustainability targets, the global production of green hydrogen is forecast to rise annually, reaching 49 million tons per year by 2030 [5]. This necessitates the design of hydrogen fuel systems to store, transport and distribute hydrogen. A long-standing challenge is hydrogen embrittlement (HE), which occurs when hydrogen atoms diffuse into metals, reducing their ductility and strength, leading to premature failure [6]. Austenitic stainless steels (SS), widely used due to their good corrosion resistance and strength, are not immune to this phenomenon. Understanding the mechanisms behind HE is essential to designing the next generation of HE-resistant materials, and will necessitate the study of hydrogen-SS interactions at different length scales.
@@
HE has been extensively studied at the macro scale through mechanical testing, revealing how cracks initiate and propagate in hydrogen-exposed material [7, 8, 9, 10] and can lead to in-service failure. To explain macroscopic behavior, advances in characterization techniques have enabled the exploration of HE at sub-micron length scales, which has been dominated by electron microscopy (EM) experiments and simulations [11, 12, 13]. This combination of experimental results and theory has led to multiple proposed degradation mechanisms responsible for HE. Here, we will briefly discuss the two most prominent mechanisms. In hydrogen-enhanced localized plasticity (HELP) [7, 8, 14], hydrogen atoms in the metal lattice accumulate at and near the dislocation cores [15] and reduce the elastic stress field, known as hydrogen elastic shielding [16]. This makes it easier for dislocations to move, leading to a localization of deformation, which drives premature material failure. The second mechanism is hydrogen-enhanced decohesion (HEDE) [17, 9], which states that hydrogen atoms accumulate at locations of high triaxial stress, such as grain boundaries or crack tips. The presence of hydrogen reduces the cohesive forces between atoms, weakening the atomic bonds and causing intergranular fracture. The interplay between HELP and HEDE is multifaceted and possibly synergistic [18, 14, 19], though there are still many different opinions about the mechanisms involved in HE [14, 20, 21]. This is because hydrogen is extremely difficult to probe due to its small size, high diffusion rate and low level of interaction with electron probes.
@@
Consequently, there has been a push towards probing in situ responses to hydrogen diffusion to understand degredation mechanisms. Dislocation movement caused by hydrogen desorption has been reported by Koyama et al. [22], where they visualised micron-scale dislocation movements using electron channel contrast imaging (ECCI) on a Fe-Mn–based alloy. Huang et al. [23] revealed bow-out motions of screw dislocations in $\alpha$-Fe pillars undergoing compression tests in a gaseous hydrogen environment using a transmission electron microscope (TEM). This is an extension to the numerous TEM experiments done by the Robertson group to study hydrogen-induced dislocation movement in foil samples [24, 25, 8, 26]. However, EM techniques only probe the surface of a material, and the question of how hydrogen interacts with a single subsurface defect in situ is still experimentally challenging and primarily addressed using simulations [13]. Furthermore, hydrogen elastic shielding has only been inferred based on the observations of dislocation pileup [27], but the hydrogen-modified strain fields have not been directly measured experimentally. This information would greatly enhance our understanding of dislocation-hydrogen interactions to accelerate the development of HE-resistant alloys.
@@
Here we address these shortcomings by using in situ Bragg coherent X-ray diffraction imaging (BCDI) [28] to probe the evolution of a single dislocation and its associated strain field during hydrogen charging (Figure 1a), using a bespoke electrochemical flow cell (Supporting Text and Figure LABEL:supp-fig:flowcell). BCDI is a nondestructive technique that captures three-dimensional (3D) nanoscale images of crystal defects and strain fields [29, 30, 31, 32, 33, 34] by inverting coherent X-ray diffraction patterns [35] from a finite crystalline sample. To produce a sample with grains suitable for BCDI, we use high pressure torsion (HPT) [36] to refine the grain structure of a 316 SS disk (composition shown in Table LABEL:supp-table:composition), followed by ex situ annealing to allow some controlled grain growth and defect relaxation (Methods). The $111$ Bragg peak from an austentitic, face centered cubic (FCC) grain is measured before and during hydrogen charging at standard ambient temperature and pressure (Methods). Phase retrieval algorithms are used to recover the real space electron density and phase from the diffraction data (Methods), revealing information about the grain morphology, dislocations, and strain fields, $\varepsilon_{111}$, along the $[111]$ direction associated with the scattering vector, $\mathbf{Q}_{111}$.
@@
Figure 1: In situ hydrogen charging BCDI setup and evolution of the Bragg peak. a), The microcrystalline 316 SS disk is mounted in the electrochemical flow cell (see Supporting Information and Figure LABEL:supp-fig:flowcell). An incoming coherent X-ray beam (red) illuminates a grain within the SS disk, and a slice through the reflected $111$ Bragg peak is captured on the detector. During in situ charging, a bias is applied to the SS (working electrode). A Pt wire coil (counter electrode) sits inside the flow cell moat. The blue arrows indicate the flow of the hydrogen charging solution. b), The lattice parameter and inferred hydrogen concentration before and during the experiment, with insets showing slices through the center of the Bragg peak corresponding to different times. c), 3D Pearson correlation of the Bragg peaks. Hydrogen charging starts at 0 h.
@@ 2.1 Bragg peak evolution
Upon initiation of charging, hydrogen atoms diffuse into the SS disk, gradually permeating the measured grain and causing localized lattice expansion [37]. This is shown as an increase in the mean lattice parameter of the grain, $a_{0}$, measured by BCDI throughout hydrogen charging (Figure 1b). We interpret this to cause a homogeneous lattice strain, $e_{111}$, relative to $a_{0}$ at -3.4 h before charging. Note that $e_{111}$ differs from $\varepsilon_{111}$: the latter is relative to the average lattice parameter at time $t$, $a_{0,t}$ (Methods) [33]. The homogeneous lattice strain results in a volumetric strain due to hydrogen occupying interstitial sites in SS [38]. The volumetric strain is combined with the relaxation volume of hydrogen in austenitic SS to estimate the local hydrogen concentration (Methods), which increases during charging (Figure 1b).
@@ 2.2 Strain at the grain surface
Figure 1b shows a large increase in average $a_{0}$ with hydrogen, leading to homogeneous strain. Does this homogeneous strain drive an increase in intergranular strain that could drive plasticity, or does hydrogen lead to a predominantly volumetric strain? If there were a substantial increase of intergranular strain, we would expect the surface strain of the grain of interest to increase due to mismatch with its neighbors. Figure 2 shows the time evolution of surface strain. We refer to the surface strain, $\varepsilon_{111,\mathrm{surf.}}$, as the $\varepsilon_{111}$ value of the surface voxel of the reconstruction as determined using an amplitude threshold (Supporting Text). Note the surface strain is limited by our estimate of the 3D spatial resolution of 12 nm, determined using the phase retrieval transfer function (Supporting Text and Figure LABEL:supp-fig:PRTF).
@@
Figure 2: Evolution of the strain field on the grain surface before and during hydrogen charging. a), Different views of the grain surface, colored by $\varepsilon_{111,\ \mathrm{surf.}}$, the strain relative to the average lattice parameter indicated at the bottom. See Figure LABEL:supp-fig:all_isosurfaces and Supporting Video S1 for reconstructions from all time points. b), Histogram distribution of the surface strain for different time points during the hydrogen charging history. c), Evolution of average surface strain. The shaded region corresponds to one standard deviation.
@@
The morphology of the grain (Figure 2a) remains stable during the hydrogen charging, however, a softening of the grain edges is noticeable after 1.9 h (Figure LABEL:supp-fig:all_isosurfaces). Hydrogen is expected to accumulate at grain boundaries [39], recently shown by Chen et al. using cryogenic atom probe tomography [15]. This may explain the apparent rounding of the grain edges in Figure 2, where hydrogen accumulates. We do not observe any grain boundary mobility as noticed in palladium thin film grains undergoing hydrogen adsorption in a previous BCDI study [32].
@@
The distribution of $\varepsilon_{111,\ \mathrm{surf.}}$ in Figure 2a is shown in Figure 2b. The mean $\varepsilon_{111,\ \mathrm{surf.}}$ is slightly compressive at all times, and the distribution remains largely unchanged during hydrogen charging. This shows that hydrogen has little effect on grain-grain misfit and thus intragranular strains that could drive plasticity. Rather, hydrogen predominantly leads to a large volumetric strain compared to $\varepsilon_{111,\mathrm{surf.}}$. The mean and standard deviation of $\varepsilon_{111,\mathrm{surf.}}$ are plotted in Figure 2c. The reduction in average $\varepsilon_{111,\mathrm{surf.}}$ with increasing charging time suggests a small average compression at grain boundaries. This is consistent with the accumulation of hydrogen at grain boundaries, leading to a slight compression of near-boundary material.
@@ 2.3 Dislocation dynamics
A phase vortex in a BCDI reconstruction indicates the spatial position of a dislocation core [29], enabling the mapping of the 3D dislocation network (Methods). Figure 3 shows the dislocation network evolution during hydrogen charging. Initially, the grain shows two small dislocations close to the surface, and one large bow-shaped dislocation that extends across the grain, hereafter referred to as the “large dislocation.” The large dislocation remains unchanged before hydrogen charging and up to 1.9 h after the start of charging. To determine its Burgers vector, we assume that it is a shear loop and lies in the $\left(\bar{1}11\right)$ plane, as expected for FCC metals (we confirm the angle between the dislocation glide plane normal and $\mathbf{Q}_{111}$ is 68.9°, less than 2° from the angle between theoretical $\langle 111\rangle$ directions and within error). Thus, we determine the orientation matrix, $\mathbf{UB}$ (Supporting Text). Glide dislocations in the $\left(\bar{1}11\right)$ plane can have three different Burgers vectors $\mathbf{b}=\frac{a_{0}}{2}[110]$, $\frac{a_{0}}{2}[101]$ or $\frac{a_{0}}{2}[0\bar{1}1]$. However, $\mathbf{Q}_{111}\cdot\mathbf{b}\neq 0$ only for $\frac{a_{0}}{2}[110]$ and $\frac{a_{0}}{2}[101]$, meaning that only dislocations with these Burgers vectors will be visible in a $\mathbf{Q}_{111}$ crystal reflection [34]. By comparing the measured strain field of the large dislocation to a 3D elastic model and dislocation dynamics simulations, we determine that the large dislocation has Burgers vector $\mathbf{b}=\frac{a_{0}}{2}[110]$ (Supporting Text, Figs. LABEL:supp-fig:dislocation_model_slices-LABEL:supp-fig:dislocation_dynamics_models, and Table LABEL:supp-table:error).
Figure 3: Evolution of the dislocation network before and during hydrogen charging. a,b,c), Columns represent a different orthogonal view of the grain and the dislocation network. The top row shows the initial morphology of the grain rendered as a grey isosurface. The subsequent rows have translucent renderings of the same grain morphology (largely unchanged, see Figure 2) along with the dislocation network at different states, with glide and climb events indicated. The dislocation dynamics simulation of glide is indicated in green. Dislocation networks are colored according to the time of observation. The black arrow indicates the Burgers vector. See Supporting Video S2 for the dislocation network at each measurement time.
@@
where $\mathbf{\xi}$ is the unit dislocation line segment direction. Here we observe climb mobility to be much higher than expected for FCC materials [41]. Climb requires diffusion of self-interstitial atoms or vacancies to the dislocation line. Our results suggest hydrogen-facilitated climb, driven by the formation and migration of vacancies due to hydrogen influx [42]. This can be rationalized by the superabundant vacancies (SAV) model proposed by Fukai [43], where hydrogen reduces the formation energy of a vacancy by trapping multiple interstitial hydrogen atoms, forming vacancy–hydrogen (vac-H) clusters that are mobile even at room temperature [44]. Furthermore, hydrogen has been demonstrated to enhance metal atom diffusion [43] and lower the vacancy migration free energy barrier in FCC metals [45], which would also drive hydrogen-facilitated climb.
Our observations suggest that once there is a sufficient hydrogen concentration within the grain, there will be enough vac-H clusters and sufficient reduction in vacancy and metal migration energy such that the osmotic force can facilitate climb. The equilibrium concentration of vacancies along the length of the large dislocation depends on the spatial variation of hydrogen concentration. Based on the magnitudes of $e_{111}$ $\left(\sim 10^{-3}\right)$ compared to $\varepsilon_{111}$ $\left(\sim 10^{-4}\right)$, increasing hydrogen concentration dominates homogeneous lattice swelling. Therefore, any variations in hydrogen concentration along the dislocation, and by extension, vacancy concentration, are small by comparison. This suggests that differences in $b_{e}$, i.e. the degree of edge character along the dislocation line, will determine which parts of the dislocation climb. Figure 4 shows a magnified view of the large dislocation before and after the climb at 3.9 h. Each segment of the dislocation is colored by edge character. The edge character is low at the end of the dislocation that does not climb, making this segment less likely to climb due to a lower $F_{\mathrm{os}}$. Conversely, the segment of the dislocation that climbs also has the greatest edge character. The climb direction (magenta arrow in Figure 4) is nearly perpendicular (89.5°) to the Burgers vector, as expected. After climb, the climbing segment lies on the $\left(11\bar{1}\right)$ plane. However, since $\mathbf{b}=\frac{a_{0}}{2}[110]$ is not one of the Burgers vectors in this plane, this segment is sessile, explaining the lack of further evolution beyond 3.9 h.

### v2

@@ Abstract
David Yang Mujan Seif Guanze He Kay Song Adrien Morez Benjamin de Jager Dmytro Nykypanchuk Ross J. Harder Wonsuk Cha Edmund Tarleton Ian K. Robinson Felix Hofmann
@@
Dr. D. Nykypanchuk
Center for Functional Nanomaterials, Brookhaven National Laboratory, Upton, NY, 11973, United States
@@ 1 Introduction
Hydrogen is an attractive energy carrier that can be produced through methods such as water electrolysis or photoelectrochemical technologies using renewable resources [1]. This green hydrogen can be used to decarbonize industries that rely heavily on fossil fuels such as aviation [2], heavy transportation [3], industrial production [4], and power generation [1]. As countries strive to reach their Net Zero 2050 sustainability targets, the global production of green hydrogen is forecast to rise annually, reaching 49 million tons per year by 2030 [5]. This necessitates the design of hydrogen fuel systems to store, transport and distribute hydrogen. A long-standing challenge is hydrogen embrittlement (HE), which occurs when hydrogen atoms diffuse into metals, reducing their ductility and strength, leading to premature failure [6]. Austenitic stainless steel (SS), widely used due to its good corrosion resistance and strength, is not immune to this phenomenon. Understanding the mechanisms behind HE is essential to designing the next generation of HE-resistant materials, and will necessitate the study of hydrogen-SS interactions at different length scales.
@@
HE has been extensively studied at the macroscale through mechanical testing, revealing how cracks initiate and propagate in hydrogen-exposed material [7, 8, 9, 10] and can lead to in-service failure. To explain macroscopic behavior, advances in characterization techniques have enabled the exploration of HE at sub-micron length scales, which has been dominated by electron microscopy (EM) experiments and simulations [11, 12, 13]. This combination of experimental results and theory has led to multiple proposed degradation mechanisms responsible for HE. Here, we will briefly discuss the two most prominent mechanisms. In hydrogen-enhanced localized plasticity (HELP) [7, 8, 14], hydrogen atoms in the metal lattice accumulate at and near the dislocation cores [15] and reduce the elastic stress field, known as hydrogen elastic shielding [16]. This makes it easier for dislocations to move, leading to a localization of deformation, which drives premature material failure. The second mechanism is hydrogen-enhanced decohesion (HEDE) [17, 9], which states that hydrogen atoms accumulate at locations of high triaxial stress, such as grain boundaries or crack tips. The presence of hydrogen reduces the cohesive forces between atoms, weakening the atomic bonds and causing intergranular fracture. The interplay between HELP and HEDE is multifaceted and possibly synergistic [18, 14, 19], though there are still many different opinions about the mechanisms involved in HE [14, 20, 21].
Explaining these complex mechanisms requires insights at the nanoscale. Hydrogen can be trapped by vacancies, interstitial sites, solutes, and precipitates, revealed using thermal desorption spectroscopy (TDS) studies [22]. Hydrogen has been reported to decrease the formation energy of vacancies, leading to an increased vacancy concentration, known as the superabundant vacancies (SAV) model [23]. Vacancies can trap multiple interstitial hydrogen atoms, revealed by positron annihilation spectroscopy (PAS) [24] and density functional theory (DFT) calculations [25]. Though less energetically favored, hydrogen also occupies octahedral interstitial sites for face-centered cubic (FCC) metals [26]. Similarly, hydrogen binds to interstitial and substitutional solutes [24], but this trapping is small compared to vacancy-solute defects [27]. Precipitates such as carbides are also strong hydrogen traps, revealed locally using cryogenic atom probe tomography [15], and some have been shown to increase hydrogen embrittlement resistance by absorbing hydrogen [28].
The ability to detect hydrogen has greatly enhanced our understanding of these mechanisms. Hydrogen microprint techniques [29], secondary ion mass spectroscopy [30] can provide a static visualization of hydrogen distribution. Scanning Kelvin probe force microscopy [31], neutron tomography [32] and electron-stimulated desorption [33] have been valuable for studying time-resolved hydrogen movement in various microstructures. Despite the availability of these techniques, hydrogen is still generally considered difficult to probe due to its small size, high diffusion rate and low level of interaction with electron probes.”
@@
Consequently, there has been a push towards probing in situ responses to hydrogen diffusion to understand degradation mechanisms. Recently, dislocation movement caused by hydrogen desorption has been reported by Koyama et al. [34], where they visualised micron-scale dislocation movements using electron channel contrast imaging (ECCI) on a Fe-Mn–based alloy. Huang et al. [35] revealed bow-out motions of screw dislocations in $\alpha$-Fe pillars undergoing compression tests in a gaseous hydrogen environment using a transmission electron microscope (TEM). This is an extension to the numerous TEM experiments done by the Robertson group to study hydrogen-induced dislocation movement in foil samples [36, 37, 38, 8, 39]. However, EM techniques only probe relatively thin subsurface volumes of a material ($\lessapprox 100$ nm) [40], and the question of how hydrogen interacts with a single dislocation within the bulk in situ is still experimentally challenging, not well understood [24, 11], and primarily addressed using simulations [13]. Furthermore, hydrogen elastic shielding has only been inferred based on the observations of dislocation pileup [41], but the hydrogen-modified strain fields have not been directly measured experimentally. This information would greatly enhance our understanding of dislocation-hydrogen interactions to accelerate the development of HE-resistant alloys.
@@
Here we address these shortcomings by using in situ Bragg coherent X-ray diffraction imaging (BCDI) [42] to probe the evolution of a single dislocation and its associated strain field during hydrogen charging (Figure 1a), using a bespoke electrochemical flow cell (Supporting Text and Figure LABEL:supp-fig:flowcell). BCDI is a nondestructive technique that captures three-dimensional (3D) nanoscale images of crystal defects and strain fields [43, 44, 45, 46, 47, 48] by inverting coherent X-ray diffraction patterns [49] from a finite crystalline sample. To produce a sample with grains suitable for BCDI, we use high pressure torsion (HPT) [50] to refine the grain structure of a 316 SS disk (composition shown in Table LABEL:supp-table:composition). Powder X-ray diffraction data shows that this leads to a deformation-induced martensitic phase transition (Figure LABEL:supp-fig:powder_XRD). The austenitic phase reappears after ex situ annealing, along with some controlled grain growth and defect relaxation (Methods). The $111$ Bragg peak from an austentitic, FCC grain is measured before and during hydrogen charging at standard ambient temperature and pressure (Methods). Phase retrieval algorithms are used to recover the real space electron density and phase from the diffraction data (Methods), revealing information about the grain morphology, dislocations, and strain fields, $\varepsilon_{111}$, along the $[111]$ direction associated with the scattering vector, $\mathbf{Q}_{111}$.
@@
Figure 1: In situ hydrogen charging BCDI setup and evolution of the Bragg peak. a), The microcrystalline 316 SS disk is mounted in the electrochemical flow cell (see Supporting Information and Figure LABEL:supp-fig:flowcell). An incoming coherent X-ray beam (red) illuminates a grain within the SS disk, and a slice through the reflected $111$ Bragg peak is captured on the detector. During in situ charging, a bias is applied to the SS (working electrode). A Pt wire coil (counter electrode) sits inside the flow cell moat. The blue arrows indicate the flow of the hydrogen charging solution. b), The lattice parameter and inferred hydrogen concentration before and during the experiment, with insets showing slices through the center of the Bragg peak corresponding to different times. The shaded areas correspond to the uncertainty associated with the Bragg peak positions. c), 3D Pearson correlation of the Bragg peaks. Hydrogen charging starts at 0 h.
@@ 2.1 Bragg peak evolution
After initiating hydrogen charging at 0 h, we observe an increase in the average lattice parameter, $a_{0}$, (Figure 1b) determined using the center position of the Bragg peak (Methods). There are several mechanisms that could explain this increase in $a_{0}$. It is favorable for hydrogen to be trapped in vacancies [24]. However, the relaxation volume for a vacancy in FCC iron is negative and even with six hydrogen atoms per vacancy only becomes slightly positive [25]. Thus, the accumulation of hydrogen-filled vacancies would not explain the substantial lattice swelling we observe. Alternatively, a Frenkel mechanism could be envisaged, where hydrogen prevents Frenkel defects from recombining, and the interstitial of the Frenkel pair (which has large positive relaxation volume) causes the observed swelling. However, this is unlikely because the formation energy of Frenkel defects is large. An energetically favorable explanation would be hydrogen occupying octahedral interstitial sites, which has been reported to cause lattice expansion [51, 26]. Thus, we conclude that the homogeneous lattice strain observed, as a result of hydrogen charging, is dominated by a volumetric strain due to hydrogen occupying interstitial sites. Here, the homogeneous lattice strain, $e_{111}$, is relative to $a_{0}$ at -3.4 h before charging. Note that $e_{111}$ differs from $\varepsilon_{111}$: the latter is relative to the average lattice parameter at time $t$, $a_{0,t}$ (Methods) [47]. The volumetric strain is combined with the relaxation volume of hydrogen in austenitic SS to estimate the local hydrogen concentration (Methods), which increases during charging (Figure 1b).
@@ 2.2 Strain at the grain surface
Figure 1b shows a large increase in average $a_{0}$ with hydrogen, leading to homogeneous strain. Does this homogeneous strain drive an increase in intergranular strain that could drive plasticity, or does hydrogen lead to a predominantly volumetric strain? If there were a substantial increase of intergranular strain, we would expect the surface strain of the grain of interest to increase due to mismatch with its neighbors. Figure 2 shows the time evolution of the heterogeneous surface strain, relative to the average lattice parameter at each time. We refer to the surface strain, $\varepsilon_{111,\mathrm{surf.}}$, as the $\varepsilon_{111}$ value of the surface voxel of the reconstruction as determined using an amplitude threshold (Supporting Text). Note the surface strain is limited by our estimate of the 3D spatial resolution of 12 nm, determined using the phase retrieval transfer function (Supporting Text and Figure LABEL:supp-fig:PRTF).
@@
Figure 2: Evolution of the heterogeneous strain field on the grain surface before and during hydrogen charging. a), Different views of the grain surface, colored by $\varepsilon_{111,\ \mathrm{surf.}}$, the strain relative to the average lattice parameter indicated at the bottom. See Figure LABEL:supp-fig:all_isosurfaces and Supporting Video S1 for reconstructions from all time points. b), Histogram distribution of the surface strain for different time points during the hydrogen charging history. c), Evolution of average surface strain. The shaded region corresponds to one standard deviation.
@@
BCDI reconstructions are very sensitive to strain $\left(\sim 10^{-4}\right)$ and highly strained regions scatter outside the detector solid angle, so they appear as missing amplitude in the reconstructions [43]. Relatively disordered grain boundaries with crystal defects in their vicinity appear rough in the reconstructions [52]. This can be seen in Figure 2a, where the morphology is stable during the hydrogen charging, however, a softening of the grain edges is noticeable after 1.9 h (Figure LABEL:supp-fig:all_isosurfaces). Grain boundaries are strong hydrogen traps ($\sim 17\ \mathrm{kJ\ mol^{-1}}$)[22], so they are expected accumulate hydrogen locally [53] as recently shown by Chen et al. using cryogenic atom probe tomography [15]. As the grain boundaries becomes better organized, here facilitated by hydrogen, lattice distortion near the boundary reduces and hence the apparent surface reconstructed from BCDI evolves. We do not observe any grain boundary mobility due to the presence of new phases, as noticed in palladium thin film grains undergoing a phase transition under hydrogen partial pressures in a previous BCDI study [46].
@@
The distribution of $\varepsilon_{111,\ \mathrm{surf.}}$ in Figure 2a is shown in Figure 2b. The mean $\varepsilon_{111,\ \mathrm{surf.}}$ is slightly compressive at all times, and the distribution remains largely unchanged during hydrogen charging. This shows that hydrogen has little effect on grain-grain misfit and thus intragranular strains that could drive plasticity. Rather, hydrogen predominantly leads to a large volumetric strain compared to $\varepsilon_{111,\mathrm{surf.}}$. The mean and standard deviation of $\varepsilon_{111,\mathrm{surf.}}$ are plotted in Figure 2c. The reduction in average $\varepsilon_{111,\mathrm{surf.}}$ with increasing charging time suggests a small average compression at grain boundaries. This is consistent with the accumulation of hydrogen at grain boundaries, leading to a slight compression of near-boundary material. Another interpretation could be the hydrogen-facilitated diffusion of chromium to the grain boundary to replenish any chromium lost during the possible formation of carbides during the annealing process, resulting in a reduction of tensile strain.
@@ 2.3 Dislocation dynamics
A phase vortex in a BCDI reconstruction indicates the spatial position of a dislocation core [43], enabling the mapping of the 3D dislocations (Methods). Figure 3 shows the dislocation evolutions during hydrogen charging. Initially, the grain shows two small dislocations close to the surface, and one large bow-shaped dislocation that extends across the grain, hereafter referred to as the “large dislocation.” The large dislocation remains unchanged before hydrogen charging and up to 1.9 h after the start of charging. To determine its Burgers vector, we assume that it is a shear loop and lies in the $\left(\bar{1}11\right)$ plane, as expected for FCC metals (we confirm the angle between the dislocation glide plane normal and $\mathbf{Q}_{111}$ is 68.9°, less than 2° from the angle between theoretical $\langle 111\rangle$ directions and within error). Thus, we determine the orientation matrix, $\mathbf{UB}$ (Supporting Text). Glide dislocations in the $\left(\bar{1}11\right)$ plane can have three different Burgers vectors $\mathbf{b}=\frac{a_{0}}{2}[110]$, $\frac{a_{0}}{2}[101]$ or $\frac{a_{0}}{2}[0\bar{1}1]$. However, $\mathbf{Q}_{111}\cdot\mathbf{b}\neq 0$ only for $\frac{a_{0}}{2}[110]$ and $\frac{a_{0}}{2}[101]$, meaning that only dislocations with these Burgers vectors will be visible in a $\mathbf{Q}_{111}$ crystal reflection [48]. By comparing the measured strain field of the large dislocation to a 3D elastic model and dislocation dynamics simulations, we determine that the large dislocation has Burgers vector $\mathbf{b}=\frac{a_{0}}{2}[110]$ (Supporting Text, Figs. LABEL:supp-fig:dislocation_model_slices-LABEL:supp-fig:dislocation_dynamics_models, and Table LABEL:supp-table:error).
Figure 3: Evolution of the dislocations before and during hydrogen charging. a,b,c), Columns represent a different orthogonal view of the grain and the dislocations. The top row shows the initial morphology of the grain rendered as a grey isosurface. The subsequent rows have translucent renderings of the same grain morphology (largely unchanged, see Figure 2) along with the dislocations at different states, with glide and climb events indicated. The dislocation dynamics simulation of glide is indicated in green. Dislocations are colored according to the time of observation. The black arrow indicates the Burgers vector. See Supporting Video S2 for the dislocations at each measurement time.
@@
where $\mathbf{\xi}$ is the unit dislocation line segment direction. Here we observe climb mobility to be much higher than expected for FCC materials [55]. Climb requires diffusion of self-interstitial atoms or vacancies to the dislocation line. Our results suggest hydrogen-facilitated climb, driven by the formation and migration of vacancies as suggested by the SAV model [23, 56]. Furthermore, hydrogen has been demonstrated to enhance metal atom diffusion [23] and lower the vacancy migration free energy barrier in FCC metals [57], which would also drive hydrogen-facilitated climb.
Our observations suggest that once there is a sufficient hydrogen concentration within the grain, there will be enough vacancies and sufficient reduction in vacancy and metal migration energy such that the osmotic force can facilitate climb. The equilibrium concentration of vacancies along the length of the large dislocation depends on the spatial variation of hydrogen concentration. Based on the magnitudes of $e_{111}$ $\left(\sim 10^{-3}\right)$ compared to $\varepsilon_{111}$ $\left(\sim 10^{-4}\right)$, increasing hydrogen concentration dominates homogeneous lattice swelling. Therefore, any variations in hydrogen concentration along the dislocation, and by extension, vacancy concentration, are small by comparison. This suggests that differences in $b_{e}$, i.e. the degree of edge character along the dislocation line, will determine which parts of the dislocation climb. Figure 4 shows a magnified view of the large dislocation before and after the climb at 3.9 h. Each segment of the dislocation is colored by edge character. The edge character is low at the end of the dislocation that does not climb, making this segment less likely to climb due to a lower $F_{\mathrm{os}}$. Conversely, the segment of the dislocation that climbs also has the greatest edge character. The climb direction (magenta arrow in Figure 4) is nearly perpendicular (89.5°) to the Burgers vector, as expected. After climb, the climbing segment lies on the $\left(11\bar{1}\right)$ plane. However, since $\mathbf{b}=\frac{a_{0}}{2}[110]$ is not one of the Burgers vectors in this plane, this segment is sessile, explaining the lack of further evolution beyond 3.9 h.
Dislocation climb is surprising since dislocation cross-slip is more energetically favored. Dislocation cross-slip requires dislocation screw character. Here, the Burgers vector is $\mathbf{b}=\frac{a_{0}}{2}[110]$, which is not parallel to the dislocation line segment that unpins, thus the segment cannot cross-slip from the $\left(\bar{1}11\right)$ plane to the $\left(11\bar{1}\right)$ plane (Figure 4). Unfortunately, we do not know what the dislocation looks like between 3.2 h and 3.9 h, so we cannot rule out cross-slip based on the absence of screw character alone. However, cross-slip also requires high stacking fault energy, which is known to be reduced by hydrogen in austenitic SS [58]. It also requires applied stress, which is minimal based on Figure 2. Thus, we propose that vacancy-assisted climb is most plausible in the hydrogen environment.